A particle of mass 15 kg has an intitial velocity vi= i cap -2 j cap m/s. It collides with another body and the impact time is 0.1 sec, resulting in a velocity vf= 6 i cap + 4 j cap + 5 k cap m/s after impact. The average force of impact on particle is????....kindly answer as soon as possible..u can also write only the answer 

Grade:11

1 Answers

Mahima Kanawat
5 years ago
Dear student
F = dP / dt
F = m [V- Vi] /t = 15 [(6i + 4j +5k) - (1i - 2j)] /0.1
   =15 [ 5i +6j +5k] /0.1
   =150 (86)1/2 =150*9.27=1391 J
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT